Temp software watchdog now operational for LO1 and the remaining optics!
Koji helped me understand how to write to switches and we tried for a while to only turnoff the output switch of the filters instead of the writing a zero that resets everything in the filter.
Eventually, I was able to move this effort foward by realising that I can pass the control trigger along multiple records using the forwarding option 'FLNK'. When I added this field to the trigger block, record(dfanout,"C1:SUS-LO1_PUSH_ALL"), and subsequent calculation blocks, record(calcout,"C1:SUS-LO1_COILSWa") to record(calcout,"C1:SUS-LO1_COILSWd"), everything started working right.
After some work on the reference database file, we now have a template for temporary watchdog implementation for LO1 located here "/cvs/cds/caltech/target/c1susaux/C1_SUS-AUX_LO1.db".
Basically, what I have done is swap the EPICS asyn analog input readout for the COIL and OSEM to accessible medm channels, then write out watchdog enable/disable to coil filter SW2 switch. Everything else in the file remains the same. I am worried about some of the conversions but the only way to know more is to see the output on the medm screen.
To test, I restarted c1su2 but this did not make the LO1 database available, so I am guessing that we also need to restart the c1sus, which can be done tomorrow.